Recently some electronic modules in the middle crate on the top level of LHRS are not working correctly. I tried to send the shower-sum signal to two different discriminators and the output are shown in figure 1 and 2. The analog signal is OK (I checked on the oscilloscope) but the discriminator outputs look funny (note the scale on the scope and how the position of ground is off in the photos. In case it's not clear in figure 2, the horizontal scale is 400 us there and the vertical is 500 mV).
Then I looked at the output from a vme pulser generator (figure 3). Some ripples can be seen there. Figure 4 shows the output from the EDTM module. I don't see any pulses but the ground is shifted a lot.
These modules have been working like that since Aug. 7. I don't know whether it has to do with the cabling for DVCS calorimeter on the first level on LHRS.